# Beyond Linear Probes: Dynamic Safety Monitoring for Language Models

### v2

@@ 1 Introduction
However, despite widespread benefits to model capabilities, dynamic computation (Han et al., 2021) for AI safety remains nascent. This is particularly true in the domain of LLM safety monitors, trained to detect harmful requests (Han et al., 2024), or problematic model behavior (Goldowsky-Dill et al., 2025; MacDiarmid et al., 2024; Chaudhary & Barez, 2025). Popular monitoring techniques include LLM-as-judges of natural language on the one hand (Inan et al., 2023; Zeng et al., 2025), and cheap linear probes in activation-space on the other (Alain & Bengio, 2017). In both cases, we argue that current approaches are inflexible. Considering that most requests are benign, dedicated LLMs have an excessively large minimum cost as always-on monitors, while activation probes provide only the most basic, static guardrails. Whilst recent work proposes to chain these two existing approaches (McKenzie et al., 2025; Cunningham et al., 2025), requiring large external LLMs that need finetuning/prompting limits their flexibility. In contrast, a single activation monitor that scales dynamically offers two key benefits:
@@
In this paper, we propose truncated polynomial classifiers (TPCs) to achieve these two properties–refining linear probes’ decision boundary by modelling rich higher-order interactions. Specifically, we show how a degree-$N$ polynomial can be trained and evaluated progressively, yielding $N$ nested submodels. Once trained, a single TPC provides dynamic defense across a range of compute budgets, through truncated evaluation. Evaluating higher-order terms provides stronger guardrails when needed, naturally generalizing the familiar linear probe (as illustrated in Fig. 1).
@@
  • We demonstrate two complementary evaluation modes of TPCs; user-driven evaluation to meet safety budgets and token-driven compute, conditional on input ambiguity.
@@ 2 Related work
Safety training of LLMs is a standard technique for preventing models from responding to problematic requests, either via post-training (Ouyang et al., 2022; Haider et al., 2024; Yuan et al., 2025; Bai et al., 2022) or during pre-training itself (O’Brien et al., 2025; Chen et al., ). Unfortunately, many models still remain vulnerable to attacks and jailbreaks (Hughes et al., 2024; Anil et al., 2024), underscoring the importance of post-training safety guardrails. One common strategy to achieve this is to use standalone LLMs trained as safety classifiers (Weng et al., ; Inan et al., 2023; Han et al., 2024; Zeng et al., 2025), leveraging LLMs’ ability to generalize to identifying novel categories of harmful inputs/responses (Inan et al., 2023). However, whilst LLMs-as-monitors are powerful, they bring significant computational cost on top of every request (Li et al., 2025), which can be prohibitively expensive for always-on monitoring.
@@ 3.1 Preliminaries
We denote matrices (vectors) using uppercase (lowercase) bold letters, e.g., $\bm{X}$ ($\bm{x}$), scalars in lowercase, e.g., $x$, and higher-order tensors in calligraphic letters, e.g., $\bm{\mathcal{{X}}}$. An element of an $N^{\text{th}}$-order tensor $\bm{\mathcal{{X}}}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ is indexed by $N$ indices, written as $\bm{\mathcal{{X}}}(i_{1},i_{2},\ldots,i_{N})\triangleq x_{i_{1}i_{2}\ldots i_{N}}\in\mathbb{R}$. We use square brackets to group weights related to the $k$-th order term in a polynomial, e.g., $w^{[k]}$. Finally, for multiple summations sharing the same upper-bound, we use the shorthand $\sum_{d_{1},d_{2},\ldots,d_{N}}^{D}$ to denote the nested summation $\sum_{d_{1}=1}^{D}\sum_{d_{2}=1}^{D}\ldots\sum_{d_{N}=1}^{D}$.
@@
We are given a dataset of $I\in\mathbb{N}$ prompts, labeled at the sequence-level as either harmful or harmless. For each input prompt, an LLM produces a $D$-dimensional residual stream representation (for each of the $T$ tokens) at a particular layer, which we denote with $\bm{Z}\in\mathbb{R}^{T\times D}$. Throughout the paper, we use single vector-valued representations of all tokens in a prompt via mean pooling with: $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{z}_{t}^{(i)}\in\mathbb{R}^{D}$. Thus, the dataset of all $I$ intermediate activations and their labels are denoted with $\mathcal{D}=\{\bm{z}^{(i)},y^{(i)}\}_{i=1}^{I}$, with each $y^{(i)}\!\in\!\{0,1\}$. For brevity, we drop the superscript indexing into a specific example in the dataset unless necessary.
@@ 3.2 Truncated polynomial classifiers
Consider a degree $N$ polynomial (Chrysos et al., 2020; Dubey et al., 2022) in the LLMs’ activation vector $\bm{z}\in\mathbb{R}^{D}$. Using the notation introduced above, we define the truncated polynomial classifier (TPC) with only its first $n\leq N$ terms as:
@@ 3.2.2 Cascading defense
Similar to recent work (McKenzie et al., 2025; Cunningham et al., 2025), we first evaluate the linear probe $y=P^{[N]}_{:1}(\bm{z})=w^{[0]}+{\bm{z}^{\top}\bm{w}^{[1]}}\in\mathbb{R}$. We then add additional higher-order terms only if the prediction falls outside a confidence threshold $\sigma(y)\notin(\tau,1-\tau)$. This cascading polynomial defense is described in Algorithm 1.
@@
4:  $y\leftarrow y+P^{[N]}_{n}(\bm{z})$ $\triangleright$ Add the $n^{\text{th}}$-order interactions
5:  if $\sigma(y)\;\notin\;(\tau,1-\tau)$ then
6:   return $y$ $\triangleright$ Early-exit with confident prediction from truncated $P^{[N]}_{:n}(\bm{z})$
7:return $y$ $\triangleright$ Otherwise return full polynomial’s prediction
@@ 3.2.3 Exploiting symmetry in model form
where a single factor matrix $\bm{U}^{[k]}\in\mathbb{R}^{R\times D}$ and coefficient vector $\bm{\lambda}^{[k]}\in\mathbb{R}^{R}$ form each degree $k$’s weights. The symmetric factorization ties weights for all permutations of the same neurons to remove redundant parameters modeling the same monomial. Whilst the regular CP decomposition reduces parameter count over Eq. 2, it still models repeated terms through multiple factor matrices. Plugging the symmetric weights in Eq. 4 into Eq. 2 yields the final truncated forward pass:
@@
with a set of learnable parameters $\bm{\theta}^{[k]}=\{\bm{\lambda}^{[k]}\in\mathbb{R}^{R},\,\,\bm{U}^{[k]}\in\mathbb{R}^{R\times D}\}$ for each degree $k>1$. Please see Appendix D for theoretical and empirical computational costs, and further discussion.
@@ 4.1 Scaling safety with test-time compute
To demonstrate the performance of TPCs on models with a range of existing guardrails, we experiment with four different LLMs of three kinds; (1) instruction-tuned model gemma-3-27b-it (Gemma Team et al., 2025), (2) non-chat base models Qwen3-30B-A3B-Base (Team, 2025) and llama-3.2-3B (Dubey et al., 2024), and (3) recent reasoning model gpt-oss-20b (Agarwal et al., 2025).
@@
Figure 3: Cascading defense: following Algorithm 1, inputs are propagated to higher-order terms only if the classification at the previous degree is uncertain (dictated by $\tau$). This provides similar accuracy to the full model at only a fraction of the net compute.
@@ 4.1.1 Results & discussion
As described in Section 3.1, we extract single vector-valued representations $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{z}_{t}^{(i)}\in\mathbb{R}^{D}$ of each prompt from the residual stream at layer $L$, mean-pooled over the token dimension. We then train a single $N=5$ degree polynomial with CP rank $R=64$ for all models. We train all models $5$ times with different random seeds.
We compute the F1 scores for every $n=\{1,\ldots,5\}$ truncated submodel $P^{[5]}_{:n}(\bm{z})$, taking the mean across the set of classifiers trained with different random seeds. We plot the results when increasing $n$ at inference-time on the WildGuardMix dataset in Fig. 2, where we find TPCs compete with or exceed the performance of the black-box EE-MLPs and MLP models alike.
Fig. 9 in the Appendix provides a further breakdown of these results by subcategory–here, we show that the full $5^{\text{th}}$-order polynomial brings up to $\textbf{10}\%$ accuracy over the linear probe for certain subcategories of harm, and up to 6% improvement over the EE-MLP in the high-compute regime. Our full results for all models and datasets are included in Appendix F.
Ultimately, TPCs provide a flexible way to extend the familiar linear probe–trading more compute for stronger guardrails at test-time, or recovering the lightweight probe exactly by evaluating the truncation $P^{[N]}_{:1}$. Please see the appendix for further ablations, including for rank (Section G.3), maximum degree (Section G.4), and computational costs (Appendix E).
@@ 4.2 Cascading defense
We next turn to demonstrate the second complementary inference-time evaluation strategy, using token-driven amounts of compute. As described in Algorithm 1, we propagate each input token to higher-order terms of the trained TPC only if the previous sub-classifier is uncertain, similar to the early-exit strategy (Teerapittayanon et al., 2016) for deep neural networks.
@@ 4.4 Pairwise feature attribution
$c_{ij}=\left(w^{[2]}_{ij}+w^{[2]}_{ji}\right)z_{i}z_{j}=\bigg(2\cdot\sum_{r=1}^{R}\lambda^{[2]}_{r}u^{[2]}_{ri}u^{[2]}_{rj}\bigg)z_{i}z_{j},$ (6)
@@ 5 Conclusion
Our experiments show impressive performance in generalizing linear probes for dynamic monitoring on large-scale safety datasets. However, we have not explored how TPCs perform in the small data regime–we anticipate stronger regularization may be needed in this setting to prevent overfitting of both TPCs and non-linear EE-MLPs probes alike. Secondly, whilst the TPC model provides built-in, mechanistically faithful explanations of exactly how much neuron combinations alter the classifier logits, the feature combinations in Section 4.4 are dense, and lack obvious legibility to humans. We are excited about future work that may use the interpretable architecture of TPCs in more interpretable bases–for example, polynomial expansions of SAE features (Tillman & Mossing, 2025; Bricken et al., 2024), and/or imposing sparsity constraints for learning interactions between only the most salient few neurons.
@@ Acknowledgments
This work was supported in part by the Technical AI Governance Initiative Fellowship at Oxford. AB acknowledges the 2025 UK AISI Systemic Safety Grant. PT acknowledges the UKRI Turing AI Fellowship (EP/W002981/1).

### v3

@@ 1 Introduction
However, despite widespread benefits to model capabilities, dynamic computation (Han et al., 2021) for AI safety remains nascent. This is particularly true in the domain of LLM safety monitors, trained to detect harmful requests (Han et al., 2024), or problematic model behavior (Goldowsky-Dill et al., 2025; MacDiarmid et al., 2024; Chaudhary & Barez, 2025). Popular monitoring techniques include LLM-as-judges of natural language on the one hand (Inan et al., 2023; Zeng et al., 2025), and cheap linear probes in activation-space on the other (Alain & Bengio, 2017). In both cases, we argue that current approaches are inflexible. Considering that most requests are benign, dedicated LLMs have an excessively large minimum cost as always-on monitors, while activation probes provide only the most basic, static guardrails. Whilst recent work proposes to chain the two existing approaches (McKenzie et al., 2025; Cunningham et al., 2025), requiring large external LLMs that need finetuning/prompting limits their flexibility. In contrast, activation monitors that scale dynamically offer two key benefits:
@@
In this paper, we propose truncated polynomial classifiers (TPCs) to achieve these two properties–refining linear probes’ decision boundary by modelling rich higher-order interactions. Specifically, we show how a degree-$N$ polynomial can be trained and evaluated progressively, yielding $N$ nested submodels. Once trained, a single TPC provides dynamic defense across a range of compute budgets, through truncated evaluation. Evaluating higher-order terms provides stronger guardrails when needed, naturally generalizing the familiar linear probe (as illustrated in Fig. 1 with a contrived example).
@@
  • We demonstrate two complementary evaluation modes of TPCs; user-driven evaluation to meet safety budgets and input-driven compute, conditional on input ambiguity.
@@ 2 Related work
Safety training of LLMs is a standard technique for preventing models from responding to problematic requests, either via post-training (Ouyang et al., 2022; Haider et al., 2024; Yuan et al., 2025; Bai et al., 2022) or during pre-training itself (O’Brien et al., 2025; Chen et al., 2025). Unfortunately, many models still remain vulnerable to attacks and jailbreaks (Hughes et al., 2024; Anil et al., 2024), underscoring the need for additional safety guardrails. One common strategy to achieve this is to use standalone LLMs trained as safety classifiers (Weng et al., 2023; Inan et al., 2023; Han et al., 2024; Zeng et al., 2025), leveraging LLMs’ ability to generalize to identifying novel categories of harmful inputs/responses (Inan et al., 2023). However, whilst LLMs-as-monitors are powerful, they bring significant computational cost on top of every request (Li et al., 2025), which can be prohibitively expensive for always-on monitoring.
@@ 3.1 Preliminaries
We denote matrices (vectors) using uppercase (lowercase) bold letters, e.g., $\bm{X}$ ($\bm{x}$), scalars in lowercase, e.g., $x$, and higher-order tensors in calligraphic letters, e.g., $\bm{\mathcal{{X}}}$. An element of an $N^{\text{th}}$-order tensor $\bm{\mathcal{{X}}}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ is indexed by $N$ indices, written as $\bm{\mathcal{{X}}}(i_{1},i_{2},\ldots,i_{N})\triangleq x_{i_{1}i_{2}\ldots i_{N}}\in\mathbb{R}$. We use square brackets to group weights related to the $k$-th order term in a polynomial, e.g., $w^{[k]}$, Finally, for multiple summations sharing the same upper-bound, we use the shorthand $\sum_{d_{1},d_{2},\ldots,d_{N}}^{D}$ to denote the nested summation $\sum_{d_{1}=1}^{D}\sum_{d_{2}=1}^{D}\ldots\sum_{d_{N}=1}^{D}$.
@@
We are given a dataset of $I\in\mathbb{N}$ prompts, labeled at the sequence-level as either harmful or harmless. For each input prompt $i$, an LLM produces a $D$-dimensional residual stream representation (for each of the $T$ tokens) at a particular layer, which we denote with $\bm{H}^{(i)}\in\mathbb{R}^{D\times T}$. Throughout the paper, we use single vector-valued representations of all tokens in a prompt via mean pooling with: $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{h}_{t}^{(i)}\in\mathbb{R}^{D}$. Thus, the dataset of all $I$ intermediate activations and their labels are denoted with $\mathcal{D}=\{\bm{z}^{(i)},y^{(i)}\}_{i=1}^{I}$, with each $y^{(i)}\!\in\!\{0,1\}$. For brevity, we drop the superscript indexing into a specific example in the dataset unless necessary.
@@ 3.2 Truncated polynomial classifiers
Consider a degree $N$ polynomial (Chrysos et al., 2020; Dubey et al., 2022) in the LLMs’ activation vector $\bm{z}\in\mathbb{R}^{D}$. Using the notation introduced above, we define the truncated polynomial classifier (TPC) up to degree $n\leq N$ as:
@@ 3.2.2 Cascading defense
Similar to recent work (McKenzie et al., 2025; Cunningham et al., 2025), we first evaluate the linear probe $s=P^{[N]}_{:1}(\bm{z})=w^{[0]}+{\bm{z}^{\top}\bm{w}^{[1]}}\in\mathbb{R}$. We then add additional higher-order terms only if the partial prediction remains uncertain, i.e., $\sigma(s)\in(\tau,1-\tau)$. This cascading polynomial defense is described in Algorithm 1.
@@
4:  $s\leftarrow s+P^{[N]}_{n}(\bm{z})$ $\triangleright$ Add the $n^{\text{th}}$-order interactions
5:  if $\sigma(s)\;\notin\;(\tau,1-\tau)$ then
6:   return $s$ $\triangleright$ Early-exit with confident prediction from truncated $P^{[N]}_{:n}(\bm{z})$
7:return $s$ $\triangleright$ Otherwise return full polynomial’s prediction
@@ 3.2.3 Exploiting symmetry in model form
where a single factor matrix $\bm{U}^{[k]}\in\mathbb{R}^{D\times R}$ and coefficient vector $\bm{\lambda}^{[k]}\in\mathbb{R}^{R}$ form each degree $k$’s weights. The symmetric factorization ties weights for all permutations of the same neurons to remove redundant parameters modeling the same monomial. Whilst the regular CP decomposition reduces parameter count over Eq. 2, it still models repeated terms through multiple factor matrices. Plugging the symmetric weights in Eq. 4 into Eq. 2 yields the final truncated forward pass:
@@
with a set of learnable parameters $\bm{\theta}^{[k]}=\{\bm{\lambda}^{[k]}\in\mathbb{R}^{R},\,\,\bm{U}^{[k]}\in\mathbb{R}^{D\times R}\}$ for each degree $k>1$. Please see Appendix D for theoretical and empirical computational costs, and further discussion.
@@ 4.1 Scaling safety with test-time compute
To demonstrate the performance of TPCs on models with a range of existing guardrails, we experiment with four different LLMs of three kinds; (1) instruction-tuned model gemma-3-27b-it (Gemma Team et al., 2025), (2) non-chat base models Qwen3-30B-A3B-Base (Qwen Team, 2025) and llama-3.2-3B (Dubey et al., 2024), and (3) recent reasoning model gpt-oss-20b (Agarwal et al., 2025).
@@
Figure 3: Cascading defense: following Algorithm 1, inputs are propagated to higher-order terms only if the classification at the previous degree is uncertain (dictated by $\tau$). This provides similar accuracy to the full model at only a fraction of the net compute (on gemma-3-27b-it at L4040).
@@ 4.1.1 Results & discussion
As described in Section 3.1, we extract single vector-valued representations $\bm{z}^{(i)}=\frac{1}{T}\sum_{t=1}^{T}\bm{h}_{t}^{(i)}\in\mathbb{R}^{D}$ of each prompt from the residual stream at layer $L$, mean-pooled over the token dimension. We then train a single $N=5$ degree polynomial with CP rank $R=64$ for all models. We train all models $5$ times with different random seeds.
Dynamic performance
We compute the F1 scores [^1] We note that we report the F1 scores throughout the paper as percentages for readability. for every $n=\{1,\ldots,5\}$ truncated submodel $P^{[5]}_{:n}(\bm{z})$, taking the mean across the set of classifiers trained with different random seeds. We plot the results when increasing $n$ at inference-time on the WildGuardMix dataset in Fig. 2, where we find TPCs compete with or exceed the performance of the black-box EE-MLPs and MLP models alike. Fig. 10 in the Appendix provides a further breakdown of these results by subcategory for gemma-3-27b-it at layer $40$. Our full results for all models and datasets are included in Appendix F, displayed as line graphs to visualize performance as a function of test-time compute.
Static performance
In addition to our main comparisons of performance with dynamic evaluation, we also take the full results from the line graphs in Section F.1 across 4 models, 2 layers, and 2 datasets (for $R=64$), and report the test set performance. We tabulate results at full depth on the layers with the best F1 score on the validation sets. Whilst this paper is primarily interested in how models perform dynamically, these results provide a complementary view of model performance at full static evaluation. These are shown in Table 1(b) for the WildGuardMix and BeaverTails datasets respectively. We observe that TPCs outperform both MLP probe variants on the challenging WildGuardMix test set, across all models considered. For BeaverTails, whilst its test set performance is slightly behind EE-MLPs for $3/4$ base models, TPCs’ max performance on the validation set remains the highest in $3/4$ comparisons.
Ultimately, TPCs provide a flexible way to extend the familiar linear probe–trading more compute for stronger guardrails at test-time, or recovering the lightweight probe exactly by evaluating the truncation $P^{[N]}_{:1}$. Please see the appendix for further ablations, including for rank (Section G.3), maximum degree (Section G.4), and computational costs (Appendix E). Please also find initial results on cross-dataset evaluation (Section F.2), more thorough multi-layer sweeps (Section F.4), and comparisons to LLM-as-monitors (Section F.3).
Table 1: Static evaluation: F1 scores at layers with best validation set performance (from full depth predictions). Results are the mean over 55 random seeds. Dynamic results found in Section F.1.
gemma-3-27B-it Qwen3-30b-A3B-Base gpt-oss-20b Llama-3.2-3B
layers: [32, 40] layers: [32, 40] layers: [16, 20] layers: [16, 20]
Method Layer Val F1 Test F1 Layer Val F1 Test F1 Layer Val F1 Test F1 Layer Val F1 Test F1
Linear probe 32 97.83 88.03 32 95.77 85.53 16 97.37 86.70 16 95.10 83.24
Bilinear probe 32 98.10 88.79 32 97.10 84.87 16 97.50 87.13 16 96.70 84.78
MLP 32 98.31 88.49 32 97.57 85.48 16 98.21 87.86 16 97.12 83.77
EE-MLP (5th exit) 32 98.22 88.39 32 97.52 85.24 16 98.08 87.31 16 96.92 83.84
TPC (5th order) 32 98.34 88.86 32 97.62 85.57 16 98.13 88.05 16 97.18 84.48
(a) WildGuardMix (Han et al., 2024)
gemma-3-27B-it Qwen3-30b-A3B-Base gpt-oss-20b Llama-3.2-3B
layers: [32, 40] layers: [32, 40] layers: [16, 20] layers: [16, 20]
Method Layer Val F1 Test F1 Layer Val F1 Test F1 Layer Val F1 Test F1 Layer Val F1 Test F1
Linear probe 40 82.46 82.28 32 79.83 80.21 16 81.13 80.86 16 80.62 80.84
Bilinear probe 40 83.48 82.97 32 83.45 82.98 16 83.25 82.91 20 83.36 82.92
MLP 40 83.38 82.89 32 83.48 82.96 20 83.35 82.83 16 83.42 83.01
EE-MLP (5th exit) 32 83.52 83.13 32 83.38 82.92 16 83.46 83.14 20 83.57 83.12
TPC (5th order) 40 83.55 83.08 32 83.56 83.00 20 83.48 82.91 20 83.52 83.05
(b) BeaverTails (Ji et al., 2023)
@@ 4.2 Cascading defense
We next turn to demonstrate the second complementary inference-time evaluation strategy, using input-driven amounts of compute. As described in Algorithm 1, we propagate each input to higher-order terms of the trained TPC only if the previous sub-classifier is uncertain, similar to the early-exit strategy (Teerapittayanon et al., 2016) for deep neural networks and more recent 2-stage cascade classifiers (McKenzie et al., 2025; Cunningham et al., 2025).
@@ 4.4 Pairwise feature attribution
$c_{ij}=\left(w^{[2]}_{ij}+w^{[2]}_{ji}\right)z_{i}z_{j}=\bigg(2\cdot\sum_{r=1}^{R}\lambda^{[2]}_{r}u^{[2]}_{ir}u^{[2]}_{jr}\bigg)z_{i}z_{j},$ (6)
@@ 5 Conclusion
Our experiments show impressive performance in generalizing linear probes for dynamic monitoring on large-scale safety datasets. However, we have not explored how TPCs perform in the small data regime–we anticipate stronger regularization may be needed in this setting to prevent overfitting of both TPCs and non-linear EE-MLPs probes alike. Secondly, whilst the TPC model provides built-in, mechanistically faithful explanations of exactly how much neuron combinations alter the classifier logits, the feature combinations in Section 4.4 are dense, and lack obvious legibility to humans. We are excited about future work that may use the interpretable architecture of TPCs in more interpretable bases–for example, polynomial expansions of SAE features (Tillman & Mossing, 2025; Bricken et al., 2024), and/or imposing sparsity constraints for learning interactions between only the most salient few neurons. We note how both TPCs and MLP baselines often fail to yield monotonically increasing performance with additional test-time compute, and all activation monitors require search for an appropriate choice of layer. We believe future work exploring more sophisticated progressive training strategies, and multi-layer probes/ensembling techniques are promising objects of future study to address such drawbacks.
@@ Acknowledgments
This work was supported in part by the Technical AI Governance Initiative Fellowship at Oxford. AB would like to acknowledge the Systematic Safety grant by UK AISI and EPSRC. PT would like to acknowledge the support of UKRI grant Turing AI Fellowship (EP/W002981/1). AB and PT are also affiliated with the Institute for Decentralized AI, which is supported by an AI Safety Fund grant.
6 Ethics statement
The goal of this work is to design better guardrails for monitoring LLMs for safer AI. We therefore feel that the work does not raise obvious or direct ethical concerns. That said, we acknowledge that the more capable model form of TPCs may inadvertently contribute to advancing AI capabilities as a side effect.
7 Reproducibility statement
To ensure the results are reproducible, we include our codebase at https://github.com/james-oldfield/tpc. Furthermore, all training details and hyperparameter sweeps used are detailed in Appendix C. Finally, a simple implementation of TPCs in PyTorch-like pseudocode is given in LABEL:lst:poly-forward.